OpenStudy (anonymous):

WILL FAN MEDAL. NEED HELP ASAP. (whatever that means)? Greg has a piece of rope that is 18 ft long. He wants to cut it into two pieces so that the longer piece is 2 ft less than 3 times the length of the shorter piece. How long will each piece of rope be? Shorter piece: ft Longer piece: ft

OpenStudy (secret-ninja):

You have to create an equation to help you. This could work: 18 ÷ 2 = 3x - 2

OpenStudy (secret-ninja):

Sorry, all of those random question marks were supposed to be a "divide by" sign.

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

18/2=3x-2?

OpenStudy (secret-ninja):

By the way, "ASAP" means "as soon as possible". :)

OpenStudy (anonymous):

oooo... lol

OpenStudy (secret-ninja):

Yep exactly. Now solve one side. 18/ 2 is 9. Find a number that can be plugged in for x.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

working on it now

OpenStudy (secret-ninja):

That will also be greater than (>) because he wants a long one and a short one.

OpenStudy (anonymous):

3x+7 ??? soo 3 is x

OpenStudy (anonymous):

yes medal plz

OpenStudy (anonymous):

The shorter piece is 3 the longer piece is ?

OpenStudy (secret-ninja):

So if 3 is x, plug it in and work it out. 18/2=3x-2 9 = 3 x 3 -2 9= 9 - 2 9 = 7 The shorter piece will be 7 feet long and the longer piece will be 9 feet long.

OpenStudy (secret-ninja):

Do you understand?

OpenStudy (anonymous):

stupid me... lol. I apologize for wasting your time. Will fan metal. and u fan metal to that other dud who wanted one?

OpenStudy (anonymous):

oh yes i understand. I knew that I just wasn't thinking.

OpenStudy (anonymous):

plz I just started

OpenStudy (secret-ninja):

Okie dokie, don't worry, you are not stupid. We all come to the time when we don't know something and have to lean it.

OpenStudy (secret-ninja):

And you didn't waste my time. Tag me on any other question you have, 'kay? Good luck. :)

OpenStudy (anonymous):

kay

OpenStudy (anonymous):

rasheed owl medal plz! I just started my account!

OpenStudy (secret-ninja):

Alex, you have to earn medals. Be happy I actually gave you one, and fanned you. I normally do NOT do that.

OpenStudy (anonymous):

YOU DID THANK YOU! lol I didn't even realize!

OpenStudy (anonymous):

Not to be rude. but your answer was incorrect, but i,m still going to be needing your help on questions... I'm still gonna tag you in my question. I got MANY MANY MANY answers incorrect before so I'm not gonna doubt you on other questions. Kay.

OpenStudy (anonymous):

Shorter-5 Longer-13 for anyone who needed to know

OpenStudy (secret-ninja):

I'm so sorry! I could have sworn i did it right... I apologize.